OpenStudy (anonymous):

write the equation y=2-4x in standard for? a.y-2=4(x-0) b.4x-y=-2 d.4x+2y=0

OpenStudy (anonymous):

Add 4x to both sides 4x + y = 2

OpenStudy (anonymous):

Just add 4x to both sides like this: y = –4x + 2 +4x + y = +4x – 4x + 2 (and since +4x – 4x = 0) 4x + y = 2